OpenStudy (anonymous):

the slope intercept form of (5/6,-4) and (3/4,-4)

OpenStudy (australopithecus):

just use the equation (y2 - y1)/(x2- x1) = m y = mx + b where m = slope b = y-intercept and (x,y) is your point

OpenStudy (anonymous):

^In addition to that, sub in the slope and either one of those points to find your "b" variable.

OpenStudy (anonymous):

Would that make y=0? Because that is incorrect.. I cannot figure out the answer.

OpenStudy (anonymous):

Why would you make y = 0? Australopithecus's solution is correct.

OpenStudy (anonymous):

(y2 - y1)/(x2- x1) = m When I plug the numbers into this equation I get 0.

OpenStudy (anonymous):

Then your m would equal 0, not your y.

OpenStudy (anonymous):

I figured that out. then my y=-4. My mistake! =)

Directrix (directrix):

Standard Form of a Linear Equation: The standard form of a linear equation is Ax + By = C, where A, B, and C are integers, A0, and A and B are not both zero. ----- If y = -4 is the correct equation, then in standard form, 0x + (1) y = -4. http://openstudy.com/study?login#/updates/4f8b4287e4b09e61bffc8324